The locus of the point which divides the double ordinates of the ellipse x^/a^+y^/b^=1 in the ratio 1:2 internally is

Grade:12

1 Answers

Arun
5 years ago
Locus will be x^2/a^2+ 9y^2/b^2=1  ,   x coordinate of any point lying in that  double ordinate will be a\cos \theta
and y cordinate will  be get from section formula . which y= (b\sin \theta)/3 . 
 
Now just square and add.
 
Regards
Arun